How much of a 90% gasoline mixture and a 75% gasoline mixture would be needed for 600 liters of an 85% gasoline mixture?
Answer by checkley77(12844)

0.90x+.75(600-x)=.85*600
.90x+450=.75x=510
.90x-.75x=510-450
.15x=60
x=60/.15
x=400 gallons of 90% gas is used.
600-400=200 gallons of 75% gas is used.
Proof:
.90*400+.75*200=.85*600
360+150=510
510=510